a radio and cycle together cost ₹1200. if the price of the former rises by 20% and then latter by 25%,then two together would cost ₹1460.find the former prices


1
Expert's answer
2022-01-21T07:34:52-0500

If r is the number of radios and c is the number of cycles before the rise, then:

r + c = 1200 and

1.2r + 1.25c = 1460,

r = 1200 - c,

"1.2\u00d7(1200 - c) +1.25c = 1460,"

0.05c = 20,

c = 400 units,

r = 1200 - 400 = 800 units.